OpenStudy (anonymous):

A triangle has sides of lengths 12, 14, and 19. Is it a right triangle? Explain.

OpenStudy (primeralph):

How do you test for right triangles?

OpenStudy (anonymous):

A. yes; 12^2 + 14^2 does not equal 19^2 B. no; 12^2 + 14^2 does not equal 19^2 C. yes; 12^2 + 14^2= 19^2 D. no; 12^2 + 14^2= 19^2

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i have no clue

OpenStudy (primeralph):

Pythagorean Theorem.

OpenStudy (anonymous):

oh god. im going to fail.

OpenStudy (primeralph):

Okay........? Sorry about that?

OpenStudy (anonymous):

ugh ill just guess.

OpenStudy (primeralph):

B.

OpenStudy (anonymous):

why'd you do that.

OpenStudy (primeralph):

Because you do it too.

OpenStudy (anonymous):

do what?